OpenStudy (anonymous):

First derivative of f(x)=3sin^2(x)

OpenStudy (anonymous):

\[f(x)=3\sin ^{2}(x)\]

OpenStudy (bahrom7893):

\[f(x) = 3(Sin(x))^2\]\[f'(x) = 2*3*Sin(x)*(Sin(x))'\]

OpenStudy (anonymous):

6sin(x)cos(x). right?

OpenStudy (bahrom7893):

Yup

OpenStudy (anonymous):

this is my original problem and that derivative isnt in the answer options!

OpenStudy (bahrom7893):

It's asking for a derivative test, not the derivative.

OpenStudy (bahrom7893):

This is tedious... ugh.. ok.. here goes..

OpenStudy (bahrom7893):

f'(x) = 0 when x = 0 \[f''(x) = -6Sin^2(x)+6Cos^2(x)=-6(Sin^2(x)-Cos^2(x))\]

OpenStudy (bahrom7893):

Now you know that f'(x) = 0 when x = 0. You need to check the signs of f'' before and after x = 0

OpenStudy (bahrom7893):

pick any negative number and evaluate that expression, then record the sign before 0. Pick any positive number and evaluate it again and record the sign after 0.

OpenStudy (anonymous):

Ok. I understand what your saying but what i dont understand Is why the first derivative isnt in the answers as "f'(x) ="

OpenStudy (anonymous):

Thats what im stuck on :/

OpenStudy (bahrom7893):

f'(x) = 0 and is near 0 are used pretty interchangeably.

OpenStudy (anonymous):

hm ok well i did what you said and plugged in numbers before and after zero & both answers were positive. by looking at my graphing calculator it sounds about right (graph is above x axis)

OpenStudy (bahrom7893):

So if they're both positive, the graph looks like this